OpenStudy (anonymous):

Use inductive reasoning to describe the pattern. Then find the next two numbers in the pattern. –9, –4, 1, 6, . . . a)add 5 to the previous term; 11, 16 b)multiply the previous term by 5; 30, 150 c)subtract 5 from the previous term; 1, –4 d)multiply the previous term by 5; 11, 150

hero (hero):

Hint: -9 + 5 = -4 -4 + 5 = 1 1 + 5 = 6 6 + 5 = ?

OpenStudy (anonymous):

@hero sorry that's why i deleted it =)